# Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

### 21 Dec 2023 v3

@@ 2.2.2 Vision foundation models (VFMs)
With the emergence of vision FMs, there is a paradigm shift towards more versatile and general-purpose models that can tackle multiple vision tasks simultaneously. These FMs leverage deep learning techniques, diverse, and large-scale training data to offer a unified framework for various computer vision tasks, enabling researchers to explore new frontiers in visual understanding and analysis (Bommasani et al., 2021). For example, Florence (Yuan et al., 2021), a newly introduced computer vision FM, offered an extensive range of capabilities for visual understanding. By leveraging universal visual-language representations from vast image-text datasets, Florence seamlessly adapted to various computer vision tasks including classification, retrieval, object detection, visual question answering (VQA), image captioning, video retrieval, and action recognition. Notably, Florence achieved outstanding performance in transfer learning scenarios, excelling in fine-tuning, linear probing, few-shot transfer, and zero-shot transfer for novel images and objects. Recently, the Segment Anything Model (SAM, (Kirillov et al., 2023)) was introduced by Meta AI for image segmentation. With its promptability and efficient architecture, SAM achieved impressive zero-shot performance on numerous segmentation tasks (Zhang et al., 2023d, Ma and Wang, 2023, Yang et al., 2023c, Tang et al., 2023). To promote research in FMs for computer vision, the SA-1B dataset, the largest of its kind, providing over 1 billion masks from 11 million images, was open-sourced [^1] SAM project https://segment-anything.com/. Inspired by SAM, the Segment Any Anomaly + (SAA+, (Cao et al., 2023b)) was proposed for zero-shot anomaly segmentation with hybrid prompt regularization to improve the adaptability of modern FMs. Furthermore, in Tang et al. (2023), the authors showed that SAM shows promising performance in camouflaged object detection (COD) tasks, in which objects are seamlessly integrated into their surroundings. In Yang et al. (2023c), SAM demonstrated superior performance in the poultry industry, specifically in the context of segmenting cage-free hens, outperforming SegFormer (Xie et al., 2021) and SETR (Liu et al., 2022) in both whole and part-based chicken segmentation. The Segmenting Everything Everywhere Model (SEEM, (Zou et al., 2023)) represented a significant advancement in the field of image segmentation. It sets a new benchmark for simultaneously segmenting everything everywhere in an image, including semantic, instance, and panoptic segmentation (Kirillov et al., 2019). Notably, SEEM supports a wide range of prompts, including visual, textual, and referring region prompts, which can be combined in various ways to enable versatile and interactive referring segmentation. This model represents a major breakthrough that pushes the boundaries of what is achievable in comprehensive image segmentation tasks.
@@ 2.2.3 Multimodal foundation models (MFMs)
Multimodal foundation models (MFMs) represent a new class of artificial intelligence models that leverage and integrate multiple types of data inputs, like text, video, audio, and images, which is a distinct departure from their non-multimodal counterparts that are constrained to processing a singular data type. By capitalizing on the richness and diversity of these data types, MFMs can attain a more holistic understanding of the context. This, in turn, empowers them to execute more nuanced and precise tasks, such as translating text into visual representations, identifying objects in images based on textual prompts, and creating image captions. As a result, these models can fuel a vast array of real-world applications (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023).
@@
One notable example is OpenAI’s CLIP (Contrastive Language-Image Pretraining, (Radford et al., 2021)), which has garnered significant attention for its ability to learn joint representations of images and text. It utilizes a strategy known as contrastive learning (Zhao et al., 2021, Jaiswal et al., 2020), which entails learning to associate accurate text-image pairs while dissociating inaccurate ones. Consequently, CLIP excels at tasks such as zero-shot learning, wherein it can classify new images based on textual descriptions and vice versa. Its multimodal nature renders it highly advantageous in diverse domains, such as computer vision, natural language processing, and digital content creation. Building on this foundation, OpenCLIP (Ilharco et al., 2021), a model trained on up to 2B image-text pairs, was proposed to investigate scaling laws (Kaplan et al., 2020) for CLIP (Radford et al., 2021) model with public datasets. OpenCLIP identified power scaling laws through various downstream tasks, such as zero-shot retrieval, classification, linear probing, and end-to-end fine-tuning, concluding that the training distribution plays a key role in scaling laws.
@@
However, a majority of foundation models are typically engineered for either understanding-based or generation-based tasks. To address this issue, BLIP (Li et al., 2022), a new Vision-Language Pre-training (VLP) framework, was introduced for both vision-language understanding and generation tasks. To effectively utilize the noisy web data by bootstrapping the captions, a captioner was employed to generate synthetic captions, subsequently, a filter was adopted to remove the noisy ones. BLIP achieved state-of-the-art results across various vision-language tasks, such as image-text retrieval, image captioning, and VQA. Similarly, OpenFlamingo (Alayrac et al., 2022) was trained on large-scale multimodal web corpora with arbitrarily interleaved text and images for a variety of image and video tasks via few-shot learning, such as image captioning and multiple-choice VQA (Alayrac et al., 2022). KOSMOS-1 (Huang et al., 2023), a Multimodal Large Language Model (MLLM), was subsequently introduced to process multimodal inputs, adhere to instructions, and perform in-context learning for language tasks and multimodal tasks, such as multimodal dialogue, image captioning, VQA, and image recognition with descriptions. Impressively, KOSMOS-1 demonstrated exemplary performance across various settings, such as zero-shot, few-shot, and multimodal chain-of-thought prompting, on various tasks without necessitating any gradient updates or fine-tuning. Built on previous versions, GPT-4 (OpenAI, 2023), a large-scale and multimodal model was introduced by OpenAI to accept both image and text inputs and produce text outputs.
@@
The future of multimodal foundation models is set to be exciting, with further advancements anticipated in several areas. They will likely include more sophisticated integration of various data modalities and enhanced transfer learning for broader applicability across different domains (Bommasani et al., 2021). Furthermore, efforts are expected to be directed towards minimizing data bias from different domains, improving fairness, and developing energy-efficient AI models (Bommasani et al., 2021).
@@ 2.2.4 Reinforcement learning foundation models (RLFMs)
Over the past decade, the field of decision making has undergone a significant revolution due to the emergence of reinforcement learning (RL) algorithms integrated with advanced DL/ML algorithms. These models, e.g., deep reinforcement learning (DRL) algorithms like DQN (Mnih et al., 2015) and PPO (Schulman et al., 2017), have enabled RL agents to learn complex behaviors and achieve remarkable performance in various domains, such as game playing (Mnih et al., 2013, 2015, Vinyals et al., 2017), robot control (Kober et al., 2013, Lee et al., 2020), and agricultural applications (Bu and Wang, 2019, Binas et al., 2019, Gandhi, 2022, Gautron et al., 2022). By interacting with the environment, these models learn optimal decision-making through trial-and-error. However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
There are increasing attempts to combing FMs and sequential decision making to tackle complex real-world problems with better generalization. Gato (Reed et al., 2022), a multi-modal, multi-task, and multi-embodiment generalist RL policy, was proposed by Google DeepMind with a 1.2B parameter decoder-only transformer architecture, trained on 604 distinct tasks with varying modalities, observations, and action specifications. Gato was able to play Atari, caption images, chat, and stack blocks with a real robot arm, using the same network and weights with little fine-tuning. Subsequently, AdA (Team et al., 2023) proposed by Google DeepMind with 265M parameters that learned new tasks in an extended XLand 3D environment simulation as fast as humans based on a modified transformer architecture to store more information and enable efficient training.
Leveraging world knowledge from FMs, sequential decision making can solve tasks faster and generalize better (Yang et al., 2023b). FMs for decision making are still in their infancy, still facing significant challenges including the discrepancy in data modalities, uncertainties surrounding environment and task structures, and the absence of certain components in existing FMs and decision-making paradigms. Further advancements are necessary to address these challenges and unlock the full potential of FMs in decision making (Yang et al., 2023b, Reed et al., 2022).
@@ 3 Agricultural Foundation Models (AFMs)
In this section, we begin by examining the potential for creating foundation models in agriculture, i.e., agricultural foundation models (AFMs). We then present current applications of these models in the realm of smart agriculture, followed by a discussion on emerging areas where these foundation models are most applicable.
@@ 3.1 Developing FMs for Agriculture
Figure 1 presents a comprehensive overview of the pipeline for developing (multimodal) foundation models in agriculture. This process principally encompasses stages such as data collection, dataset curation, model training, and downstream fine-tuning.
@@
The development of multimodal foundation models for agricultural applications commences with the imperative phase of data collection and curation (Bommasani et al., 2021). This process necessitates the assembly of heterogeneous data spanning a range of modalities, including visual (e.g., images and videos indicating the status of crops, pests, and disease symptoms), textual (e.g., human instructions and pest identification documents), meteorological (e.g., temperature, precipitation, and wind speed measurements), and possibly auditory data (e.g., distinctive insect or animal sounds). Given the magnitude of the data involved, meticulous annotation and curation are indispensable to discriminate meaningful patterns and correlations. Such data labeling could encompass tasks such as the classification of crop and weed species, the annotation of disease symptoms present on plant foliage, or the categorization of weather conditions.
@@
Subsequent to data curation is the model training phase. Capitalizing on contemporary advancements in machine learning and artificial intelligence, such as transformer architectures (Vaswani et al., 2017) and convolutional neural networks (Gu et al., 2018), the model is trained to process and interpret multimodal inputs. These artificial intelligence models are designed to associate visual depictions with their corresponding textual labels, comprehend textual descriptions, and generate an inclusive understanding of the agricultural ecosystem. Notably, the training of these models may be computationally demanding, which necessities the use of advanced hardware and optimization algorithms to enhance performance, which will be discussed in Section  4.2.
@@
The culmination of this process is the implementation of these trained models in downstream agricultural applications, such as plant health monitoring (Xu et al., 2022a), crop and weed management (Hu et al., 2021, Dang et al., 2023, Rai et al., 2023), fruit picking (Chu et al., 2021), and precision livestock farming (Yang et al., 2023c). For instance, in relation to plant health, the models can be utilized to identify early indicators of disease or pest infestations, thus facilitating prompt interventions to reduce crop loss. For weed management, the models can assist in distinguishing weeds from crops, enabling precision farming equipment to selectively eradicate unwanted flora. Furthermore, in the domain of precision livestock farming, these models can be harnessed to monitor livestock health and behavior, optimize feeding schedules, and predict potential diseases. In the context of downstream tasks, transfer learning or fine-tuning (Weiss et al., 2016, Zhuang et al., 2020) of the foundation models becomes paramount to adapt the generalized learning to specific agricultural applications, thereby enhancing model performance and efficiency. Through these varied applications, multimodal foundation models hold the potential to significantly transform the agricultural industry, thereby enhancing productivity, promoting sustainability, and ultimately boosting profitability.
However, constructing an agricultural foundation model from scratch presents significant challenges. Firstly, initial data collection requires a large and diverse dataset, which requires laborious efforts and potentially high costs (Lu and Young, 2020). The subsequent data labeling and curation require meticulousness and expertise. Additionally, the process of training models demands substantial computational resources, thus presenting impediments for individuals or entities with constrained resources. Lastly, fine-tuning of the model for specific agricultural applications can be challenging, requiring a balance between generalizability and specificity. A detailed discussion of these challenges will be presented in Section  4.
@@ 3.2 Agricultural applications
In this subsection, we explore the existing applications of integrating FMs discussed in Section  2 into the field of smart agriculture.
@@ 3.2.1 LFMs in agriculture
Language Foundation Models (LFMs), also referred to as Large-Scale Language Modeling (LLMs) and detailed in Section  2.2.1, have transformed the realms of design, development, and deployment by enabling human-like interactions. Researchers are now increasingly integrating these LFMs into various stages of design and development for agricultural applications, as highlighted in (Stella et al., 2023, Lu et al., 2023). Specifically, in Stella et al. (2023), the authors discussed the process of integrating Large Language Models (LLMs) into the design phase of robotic systems, as depicted in Figure 2 from Stella et al. (2023). They specifically illustrate the procedure to design a robotic gripper optimized for tomato picking. During the initial ideation phase, researchers consult LLMs, such as ChatGPT (OpenAI, 2022), to understand potential challenges and opportunities within the field. Based on this information, they select the most promising and intriguing pathways, subsequently narrowing down the design possibilities through further dialogue with the LLM. This collaborative interaction covers a wide spectrum of knowledge domains and various levels of abstraction, encompassing everything from conceptual thinking to technical implementation. Throughout this process, the human collaborator leverages the AI’s knowledge to access information beyond their personal expertise. During the subsequent phase of the design process, which is more technically focused, these general directions need to be actualized into a tangible, fully functional robot. As of now, LLMs are not capable of producing complete CAD models, assessing code, or autonomously fabricating a robot. However, recent progress in AI research has indicated that these algorithms can provide substantial assistance in executing software (Chen et al., 2021), mathematical reasoning (Wolfram, 2023), and even in the generation of shapes (Ramesh et al., 2022). Although we foresee that future AI methodologies will be capable of handling these tasks, at the moment, the technical implementation remains a team effort between AI models and humans. Finally, the human fine-tuned the code suggested by the LLM, finalized the CAD design, and oversaw the actual fabrication of the robot. Figure 2 (b) from Stella et al. (2023) showed the primary outputs generated by the LLM, as well as the real-world application of the AI-designed robotic gripper used for crop harvesting.
@@
The authors in Lu et al. (2023) have deliberated on the prospective use cases of Large Language Models (LLMs) in organizing unstructured metadata, converting metadata from one format to another, and identifying potential errors during the data collection phase. They also envision the future generation of LLMs as exceptionally powerful tools for data visualization (Bubeck et al., 2023), offering invaluable assistance to researchers in extracting meaningful insights from vast volumes of phenotypic data.

### 17 Mar 2024 v4

@@ 2.2.2 Vision foundation models (VFMs)
With the emergence of vision FMs, there is a paradigm shift towards more versatile and general-purpose models that can tackle multiple vision tasks simultaneously. These FMs leverage deep learning techniques, and diverse, and large-scale training data to offer a unified framework for various computer vision tasks, enabling researchers to explore new frontiers in visual understanding and analysis (Bommasani et al., 2021). For example, Florence (Yuan et al., 2021), a newly introduced computer vision FM, offers an extensive range of capabilities for visual understanding. By leveraging universal visual-language representations from vast image-text datasets, Florence seamlessly adapts to various computer vision tasks including classification, retrieval, object detection, visual question answering (VQA), image captioning, video retrieval, and action recognition. Notably, Florence achieves outstanding performance in transfer learning scenarios, excelling in fine-tuning, linear probing, few-shot transfer, and zero-shot transfer for novel images and objects. Recently, the Segment Anything Model (SAM, (Kirillov et al., 2023)) is introduced by Meta AI for image segmentation. With its flexible and efficient architecture, SAM achieved impressive zero-shot performance on numerous segmentation tasks (Zhang et al., 2023e, Ma and Wang, 2023, Yang et al., 2023c, Tang et al., 2023). To promote research in FMs for computer vision, the SA-1B dataset, the largest of its kind, providing over 1 billion masks from 11 million images, is open-sourced [^2] SAM project https://segment-anything.com/. Inspired by SAM, the Segment Any Anomaly + (SAA+, (Cao et al., 2023b)) is proposed for zero-shot anomaly segmentation with hybrid prompt regularization to improve the adaptability of modern FMs. Furthermore, in Tang et al. (2023), the authors show that SAM shows promising performance in camouflaged object detection (COD) tasks, in which objects are seamlessly integrated into their surroundings. In Yang et al. (2023c), SAM demonstrates superior performance in the poultry industry, specifically in the context of segmenting cage-free hens, outperforming SegFormer (Xie et al., 2021) and SETR (Liu et al., 2022) in both whole and part-based chicken segmentation. The Segmenting Everything Everywhere Model (SEEM, (Zou et al., 2023)) represents a significant advancement in the field of image segmentation. It sets a new benchmark for simultaneously segmenting everything everywhere in an image, including semantic, instance, and panoptic segmentation (Kirillov et al., 2019). Notably, SEEM supports a wide range of prompts, including visual, textual, and referring region prompts, which can be combined in various ways to enable versatile and interactive referring segmentation. This model represents a breakthrough that pushes the boundaries of what is achievable in comprehensive image segmentation tasks.
@@ 2.2.3 Multimodal foundation models (MFMs)
MFMs represent a new class of AI models that leverage and integrate multiple types of data inputs, like text, video, audio, and images, which is a distinct departure from their non-multimodal counterparts that are constrained to processing a singular data type. By capitalizing on the richness and diversity of these data types, MFMs can attain a more holistic understanding of the context. This, in turn, empowers them to execute more nuanced and precise tasks, such as translating text into visual representations, identifying objects in images based on textual prompts, and creating image captions. As a result, these models can fuel a vast array of real-world applications (Bommasani et al., 2021, Moor et al., 2023, Mai et al., 2023).
@@
One notable example is OpenAI’s CLIP (Contrastive Language-Image Pretraining, (Radford et al., 2021)), which has garnered significant attention for its ability to learn joint representations of images and text. It utilizes a strategy known as contrastive learning (Zhao et al., 2021, Jaiswal et al., 2020), which entails learning to associate accurate text-image pairs while dissociating inaccurate ones. Consequently, CLIP excels at tasks such as zero-shot learning, wherein it can classify new images based on textual descriptions and vice versa. Its multimodal nature renders it highly advantageous in diverse domains, such as computer vision, natural language processing, and digital content creation. Building on this foundation, OpenCLIP (Ilharco et al., 2021), a model trained on up to 2B image-text pairs, is proposed to investigate scaling laws (Kaplan et al., 2020) for CLIP (Radford et al., 2021) model with public datasets. OpenCLIP identifies power scaling laws through various downstream tasks, such as zero-shot retrieval, classification, linear probing, and end-to-end fine-tuning, concluding that the training distribution plays a key role in scaling laws. The new Gemini 1.5333Gemini 1.5 website: https://deepmind.google/technologies/gemini/#introduction. developed by Google delivers dramatically enhanced performance with a more efficient architecture compared to its previous version (Team et al., 2023b).
@@
However, a majority of FMs are typically engineered for either understanding-based or generation-based tasks. To address this issue, BLIP (Li et al., 2022), a new Vision-Language Pre-training (VLP) framework, is introduced for both vision-language understanding and generation tasks. To effectively utilize the noisy web data by bootstrapping the captions, a captioner is employed to generate synthetic captions, subsequently, a filter is adopted to remove the noisy ones. BLIP achieves state-of-the-art results across various vision-language tasks, such as image-text retrieval, image captioning, and VQA. Similarly, OpenFlamingo (Alayrac et al., 2022) is trained on large-scale multimodal web corpora with arbitrarily interleaved text and images for a variety of image and video tasks via few-shot learning, such as image captioning and multiple-choice VQA (Alayrac et al., 2022). KOSMOS-1 (Huang et al., 2023), a Multimodal Large Language Model (MLLM), is subsequently introduced to process multimodal inputs, adhere to instructions, and perform in-context learning for language tasks and multimodal tasks, such as multimodal dialogue, image captioning, VQA, and image recognition with descriptions. Impressively, KOSMOS-1 demonstrates exemplary performance across various settings, such as zero-shot, few-shot, and multimodal chain-of-thought prompting, on various tasks without necessitating any gradient updates or fine-tuning. Built on previous versions, GPT-4 (OpenAI, 2023), a large-scale and multimodal model is introduced by OpenAI to accept both image and text inputs and produce text outputs.
@@
The future of MFMs is set to be exciting, with further advancements anticipated in several areas. They will likely include more sophisticated integration of various data modalities and enhanced transfer learning for broader applicability across different domains (Bommasani et al., 2021). Furthermore, efforts are expected to be directed towards minimizing data bias from different domains, improving fairness, and developing energy-efficient AI models (Bommasani et al., 2021).
@@ 2.2.4 Reinforcement learning foundation models (RLFMs)
Over the past decade, the field of decision-making has undergone a significant revolution due to the emergence of RL algorithms integrated with advanced DL/ML algorithms. These models, e.g., deep RL (DRL) algorithms like DQN (Mnih et al., 2015) and PPO (Schulman et al., 2017), have enabled RL agents to learn complex behaviors and achieve remarkable performance in various domains, such as game playing (Mnih et al., 2013, 2015, Vinyals et al., 2017), robot control (Kober et al., 2013, Lee et al., 2020), and agricultural applications (Bu and Wang, 2019, Binas et al., 2019, Gandhi, 2022, Gautron et al., 2022). By interacting with the environment, these models learn optimal decision-making through trial and error. However, due to their reliance on learning tasks from scratch without comprehensive knowledge of vision, language, or other datasets, these methods often face challenges in terms of generalization and sample efficiency (Yang et al., 2023b).
There are increasing attempts to combine FMs and sequential decision-making to tackle complex real-world problems with better generalization. Gato (Reed et al., 2022), a multi-modal, multi-task, and multi-embodiment generalist RL policy, is proposed by Google DeepMind with a 1.2B parameter decoder-only transformer architecture, trained on 604 distinct tasks with varying modalities, observations, and action specifications. Gato is able to play Atari, caption images, chat, and stack blocks with a real robot arm, using the same network and weights with little fine-tuning. Subsequently, AdA (Team et al., 2023a) proposes by Google DeepMind with 265M parameters that learn new tasks in an extended and 3D environment simulation as fast as humans based on a modified transformer architecture to store more information and enable efficient training.
Leveraging world knowledge from FMs, sequential decision making can solve tasks faster and generalize better (Yang et al., 2023b). FMs for decision making are still in their infancy and still facing significant challenges including the discrepancy in data modalities, uncertainties surrounding environment and task structures, and the absence of certain components in existing FMs and decision-making paradigms. Further advancements are necessary to address these challenges and unlock the full potential of FMs in decision making (Yang et al., 2023b, Reed et al., 2022).
@@ 3 Agricultural Foundation Models (AFMs)
In this section, the potential for creating FMs in agriculture is first examined, i.e., AFMs. Then current applications of these models in the realm of smart agriculture are presented, followed by a discussion on emerging areas where these FMs are most applicable.
@@ 3.1 Developing FMs for Agriculture
Figure 1 presents a comprehensive overview of the pipeline for developing (multimodal) FMs in agriculture. This process principally encompasses stages such as data collection, dataset curation, model training, and downstream fine-tuning.
@@
The development of MFMs for agricultural applications commences with the imperative phase of data collection and curation (Bommasani et al., 2021). This process necessitates the assembly of heterogeneous data spanning a range of modalities, including visual (e.g., images and videos indicating the status of crops, pests, and disease symptoms), textual (e.g., human instructions and pest identification documents), meteorological (e.g., temperature, precipitation, and wind speed measurements), and possibly auditory data (e.g., distinctive insect or animal sounds). Given the magnitude of the data involved, meticulous annotation and curation are indispensable to discriminate meaningful patterns and correlations. Such data labeling could encompass tasks such as the classification of crop and weed species, the annotation of disease symptoms present on plant foliage, or the categorization of weather conditions.
@@
Subsequent to data curation is the model training phase. Capitalizing on contemporary advancements in ML and AI, such as transformer architectures (Vaswani et al., 2017) and convolutional neural networks (Gu et al., 2018), the model is trained to process and interpret multimodal inputs. These AI models are designed to associate visual depictions with their corresponding textual labels, comprehend textual descriptions, and generate an inclusive understanding of the agricultural ecosystem. Notably, the training of these models may be computationally demanding, which necessitates the use of advanced hardware and optimization algorithms to enhance performance, which will be discussed in Section  4.2.
@@
The culmination of this process is the implementation of these trained models in downstream agricultural applications, such as plant health monitoring (Xu et al., 2022b), crop and weed management (Hu et al., 2021, Dang et al., 2023, Rai et al., 2023), fruit picking (Chu et al., 2021), and precision livestock farming (Yang et al., 2023c). For instance, in relation to plant health, the models can be utilized to identify early indicators of disease or pest infestations, thus facilitating prompt interventions to reduce crop loss. For weed management, the models can assist in distinguishing weeds from crops, enabling precision farming equipment to selectively eradicate unwanted flora. Furthermore, in the domain of precision livestock farming, these models can be harnessed to monitor livestock health and behavior, optimize feeding schedules, and predict potential diseases. In the context of downstream tasks, transfer learning or fine-tuning (Weiss et al., 2016, Zhuang et al., 2020) of the FMs becomes paramount to adapt the generalized learning to specific agricultural applications, thereby enhancing model performance and efficiency. Through these varied applications, MFMs hold the potential to significantly transform the agricultural industry, thereby enhancing productivity, promoting sustainability, and ultimately boosting profitability.
@@ 3.2 Agricultural applications
In this subsection, the existing applications of integrating FMs discussed are explored in Section 2 into the field of smart agriculture.
@@ 3.2.1 LFMs in agriculture
LFMs, also referred to as LLMs and detailed in Section  2.2.1, have transformed the realms of design, development, and deployment by enabling human-like interactions. Researchers are now increasingly integrating these LFMs into various stages of design and development for agricultural applications, as highlighted in (Stella et al., 2023, Lu et al., 2023). Specifically, in Stella et al. (2023), the authors discuss the process of integrating LLMs into the design phase of robotic systems, as depicted in Figure 2 from Stella et al. (2023). They specifically illustrate the procedure to design a robotic gripper optimized for tomato picking. During the initial ideation phase, researchers consult LLMs, such as ChatGPT (OpenAI, 2022), to understand potential challenges and opportunities within the field. Based on this information, they select the most promising and intriguing pathways, subsequently narrowing down the design possibilities through further dialogue with the LLM. This collaborative interaction covers a wide spectrum of knowledge domains and various levels of abstraction, encompassing everything from conceptual thinking to technical implementation. Throughout this process, the human collaborator leverages the AI’s knowledge to access information beyond their personal expertise. During the subsequent phase of the design process, which is more technically focused, these general directions need to be actualized into a tangible, fully functional robot. As of now, LLMs are not capable of producing complete computer-aided design (CAD) models, assessing code, or autonomously fabricating a robot. However, recent progress in AI research has indicated that these algorithms can provide substantial assistance in executing software (Chen et al., 2021), mathematical reasoning (Wolfram, 2023), and even in the generation of shapes (Ramesh et al., 2022). Although future AI methodologies will be capable of handling these tasks, at the moment, the technical implementation remains a team effort between AI models and humans. Finally, the human would fine-tune the code suggested by the LLM, finalize the CAD design, and oversee the actual fabrication of the robot. Figure 2(b) in Stella et al. (2023) shows the primary outputs generated by the LLM, as well as the real-world application of the AI-designed robotic gripper used for crop harvesting.
@@
The authors in Lu et al. (2023) have deliberated on the prospective use cases of LLMs in organizing unstructured metadata, converting metadata from one format to another, and identifying potential errors during the data collection phase. They also envision the future generation of LLMs as exceptionally powerful tools for data visualization (Bubeck et al., 2023), offering invaluable assistance to researchers in extracting meaningful insights from vast volumes of phenotypic data. Furthermore, in Tzachor et al. (2023), the authors utilize GPT to revolutionize agricultural extension services by simplifying scientific knowledge and delivering personalized, location-specific, and data-driven agricultural recommendations. In Shutske (2023), Generative AI and LLMs, particularly ChatGPT, are investigated to elevate innovation and enhance agricultural safety, occupational health, and health promotion.
Moreover, researchers have leveraged LLMs for various plant science applications. For example, in Yang et al. (2024), PLLaMa, an open-source LLM based on LLaMa-2, is developed specifically for plant science-related topics. Additionally, in Kuska et al. , an AI-Chatbots for Agriculture is created to aid in interpreting decision support models in plant disease management.